OpenStudy (maheshmeghwal9):

Solve:

OpenStudy (maheshmeghwal9):

If \[2^x=\cos (y/2) \] & \[a^x=\sin y\] Then, Find \[\sin (y/2).\]

OpenStudy (maheshmeghwal9):

Please solve:)

OpenStudy (anonymous):

you know the formula : \[\sin 2u = 2\sin u \cos u\] ?

OpenStudy (maheshmeghwal9):

Ya!

OpenStudy (anonymous):

Then you have \[a^x=2\sin\frac{y}{2} \cos\frac{y}{2}\]

OpenStudy (anonymous):

a^x/2^(x+1)

OpenStudy (maheshmeghwal9):

ok !anakhoava

OpenStudy (anonymous):

\[siny=2\sin \frac y2 \cos \frac y2\]

OpenStudy (maheshmeghwal9):

Thanx a lot u all!